OpenStudy (anonymous):

integrate part of sine in interval x=[0, pi/2] function: f(x) = sin(x) Anti-Derivative: F(x) = -1/x cos(x)

OpenStudy (anonymous):

Is the anti derivative correct ?

OpenStudy (kainui):

The antiderivative of sin(x) is not -1/x cos(x). Just take the derivative of that "antiderivative" to see that the product rule will give you something with 1/x^2 *cos(x). But you're not completely wrong in guessing that. If it was sin(ax) where a is just a constant, like 2, then you would say it's -1/a*cos(ax). Take the derivative of this antiderivative to check and make sure I'm not lying to you!

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

x derived is 1......

OpenStudy (anonymous):

by x

OpenStudy (anonymous):

F(x) = (1/ 1 ) -cos(x)

OpenStudy (anonymous):

@Kainui is that the right anti- derivative? :)

OpenStudy (anonymous):

f(x) = sin(x) then F(x) = -cos(x)

OpenStudy (kainui):

That is the correct antiderivative!

OpenStudy (anonymous):

ok thanks :) so then I can take the end point - start point: \[\left[ -\cos(x)\right] \frac{ to }{ from }\]-cos(pi/2) - -cos(0) -cos(pi/2) + cos(0) -0 + 1 = 1